3.2 Show your calculation and explain how base 60 number system is used to subtract 48 minutes from 2 hour 13 minutes?


1
Expert's answer
2022-06-09T10:42:10-0400
"2h\ 13min-48min=2h+13min-48min""=1h+60min+13min-48min""=1h+73min-48min""=1h+25min=1h\ 25min"

 1 hour 25 minutes